OceanGate sells external hard drives for $250 each. Its total fixed costs are $30 million, and its variable costs per unit are $200. The corporate tax rate is 30%. If the economy is strong, the firm will sell 2 million drives, but if there is a recession, it will sell only half as many.

a. What will be the percentage decline in sales if the economy enters a recession?

b. What will be the percentage decline in profits if the economy enters a recession?

c. Comparing your answers to (a) and (b), how would you measure the operating leverage of this firm?
Answer #1

n1=2mn (Strong economy) n2=1mn(recession),

% change in sales=(250*2)-(250*1)/(250*2)

=250/500=50% decline in sales.

(Ebit)=(sales-vc)-fc

(500-400)-30=Ebit1=70mn

Ebit2=(250-200)-30=20

%change in ebit=70-20/70=0.7143 =71.43% decline in profits.

=Hence operating leverage =%change in sales/%change in ebit

=50/70=0.7143.